# A new paradigm of mortality modeling via individual vitality dynamics

### v2

@@ 3.2.3 A dynamic vitality model with age-cohort effect
    The initial vitality $V_{x,y}(0)$ for an individual born in $y$ and age $x$ at time $t=0$ follows
    $V_{x,y}(0)\sim\text{Exp}\left(\Gamma(y)\cdot a^{x}\right),$
    where $\Gamma(y)\cdot a^{x}$ is the rate parameter. We assume $a>1$ is a constant value accounting for the fact that vitality decreases as the individual ages, and $\Gamma(y)$ is governed by a stochastic process, for example, the Geometric Brownian Motion specified in Section 3.2.2.
@@
    The trend component $Y(t)$ is specified as
    $\mathop{}\!\mathrm{d}Y(t)=b(t)\frac{c(t)^{x+t}}{a^{x}}\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
    with the dynamics of $b(t)$ and $c(t)$ the same as in Sections 3.2.1 and 3.2.2.
In the above specification, $\Gamma(y)$ is designed to capture the general level of initial vitality at birth for cohort year $y$, with the term $\Gamma(y)\cdot a^{x}$ reflecting the expected vitality level when the individual reaches age $x$. On the other hand, $b(t)$ and $c(t)$ are the Gompertz parameters that capture the change in vitality depletion rate when the individual is aging at time $t$. It is immediately observed that the resulting survival probability matches with the M6 model.
It is worth noting that this specification provides more flexibility in modeling the interactions between cohorts and ages. For example, the initial vitality component can be modified to
$V_{x,y}(0)\sim\text{Exp}\left(\Gamma(y)^{x_{c}-x}\right),$
where $x_{c}$ is a large and fixed value, e.g., $x_{c}=100$. If $\Gamma(y)>1$, the cohort effect will diminish quickly as the individual ages, which reconciles with the cohort effect of the M8 model.
@@ 4 Model Applications
The proposed vitality-based modeling framework has great potential in actuarial research. In this section, we provide a number of illustrative examples to showcase the applications of a vitality mortality model. Specifically, we examine the following topics:
@@
Applying the vitality mortality model differs from existing mortality models not only in terms of technical details but also by offering new interpretations and insights that cannot be explained using existing mortality models.
@@ 4.1 Life insurance valuation
If the density function of $\tau(v)$ can be explicitly derived, then the life expectancy, annuity price, and insurance price can also be expressed explicitly in an integral form. This is the case for our proposed model if the diffusion component is not considered. When the density function of $\tau(v)$ cannot be explicitly written, an approximation method is then needed to provide accurate estimates for life expectancy and prices. For example, consider a vitality mortality model with initial vitality, trend, and diffusion (i.e., components C1-C3), then the problem becomes finding the first exit density of Brownian motion with a curved boundary. A number of studies, including Jennen and Lerche, (1981), Jennen, (1985), Skiadas, (1995), and Durbin and Williams, (1992), has provided accurate and easy-to-implement approximation methods to handle this problem. Skiadas and Skiadas, (2010) and Skiadas and Skiadas, (2014) have specifically applied approximation methods to vitality modeling. For more details on approximating the density function of the death time when C4 is absent, interested readers are referred to Appendix D.
@@ 4.2 Optimal investment and consumption strategies
One of the potential applications of vitality modeling is lifetime planning, e.g., Merton’s portfolio problem. Individuals, when making decisions about their investment portfolio, consumption, insurance purchases, medical expenses, educational investment, etc., will take into account their health level at the time. The vitality mortality model provides a natural way to frame the optimization problem. There exist a few studies that utilize the vitality-based approach. For example, Chen et al., (2022) modeled vitality as a geometric Brownian motion to study the optimal insurance decision, and Cheng et al., (2023) modeled vitality using a jump process to study the optimal medical investment. In this subsection, we provide an illustration of Merton’s portfolio consumption problem under our vitality mortality model.
@@ 4.2.1 The market and vitality processes
Following the seminal work of Merton, (1969), we construct a continuous-time model for the financial market and apply the vitality model in Section 3.1.4. Denote $(\Omega,\mathbb{F},\mathbb{P})$ as a complete probability space with a right-continuous filtration generated by a two-dimensional standard Brownian Motion $(B_{S}(t),B_{V}(t))$. An individual allocates two representative assets, one risk-free $S_{0}(t)$ and one risky $S_{1}(t)$ with the following dynamics:
@@
$\mathop{}\!\mathrm{d}S_{1}(t)$ $=(r+\theta\sigma)S_{1}(t)\mathop{}\!\mathrm{d}t+\sigma_{S}S_{1}(t)\mathop{}\!\mathrm{d}B_{S}(t),\quad S_{1}(0)>0,$
@@
Let $\pi(t)$ be the portfolio weight invested in the risky asset, and $\zeta(t)$ be the consumption amount. Then the asset $A(t)$ of an individual has the following dynamic
$\mathop{}\!\mathrm{d}A(t)=A(t)\left[r+\pi(t)\theta\sigma_{S}\right]\mathop{}\!\mathrm{d}t+\sigma\pi(t)A(t)\mathop{}\!\mathrm{d}B_{S}(t)-\zeta(t)\mathop{}\!\mathrm{d}t,\quad A(0)=A_{0}.$
We now adopt the vitality mortality model from Section 3.1.4, which recovers the Gompertz law if the initial vitality follows the Gompertz distribution; that is,
@@ 4.2.2 The optimal investment and consumption problem
where $\tau$ is the first passage time that vitality level crosses zero, $\beta>0$ is the time-preference rate, $u(\zeta)=\ln\zeta$ is the utility function (log-utility is selected for illustration), $\lambda\geq 0$ is the weight given to the utility of terminal asset (i.e., bequest), $\mathbb{E}_{t,a,v}=\mathbb{E}[\cdot|t,A(t)=a,V(t)=v]$, and the investment/consumption strategy $(\pi=(\pi_{t})_{t\geq 0},\zeta=(\zeta_{t})_{t\geq 0})\in\Pi$ belong to the admissible set $\Pi$ if $\pi$ and $\zeta$ are progressively measurable and the resulting asset process $(A_{t})_{t\geq 0}$ is non-negative for all $t$. Notice that the optimization is time-independent since $V(t)$ is stationary and thus the value function does not depend on the time $t$.
@@
As for comparison, we have Merton’s optimal portfolio and consumption strategies presented in Appendix C. The optimal investment strategy follows the same as Merton’s optimal portfolio in an infinite time horizon, and the consumption strategy also collapsed to Merton’s case (w/o mortality) if vitality is infinite.
@@ 4.2.3 Further remarks
It is not surprising to observe that the individual’s consumption depends on the vitality level, but what is interesting here is that they are not always negatively related, as one would expect. When the bequest motive is moderate or low (i.e., $\lambda<1/\beta$) a reasonable consumption profile is observed such that a higher vitality represents a long time to live with, and thus a lower consumption must be made to make sure that an individual will not run of out money before they are deceased.
However, when the bequest motive is significantly large (i.e., $\lambda>1/\beta$) decreasing the vitality actually increases the consumption, a counter-intuitive result at first glance. This is due to the fact that individuals are overly concerned about their bequest, and thus leads to an overly conservative consumption initially. Once they are aging and the time for death becomes more certain, they become more aware of their over-saving and may wish to increase their consumption a bit.
@@ 4.3 Subjective beliefs
Another interesting observation that can be reflected by our vitality model is the distinction between subjective belief in health status and subjective belief in life expectancy. Specifically, if a person correctly self-assesses as having an average health level among the whole population, they may still over- or under-estimate their life expectancy if they falsely believe that their life expectancy will also be the average among the whole population. Using the specification from Section 3.1.1 with $b=0.0001744$ and $c=1.082$ as an illustrative example, the population average life expectancy at age 60 is 17 years. However, a 60-year-old individual who believes that they have an average initial vitality level (which is one since the initial vitality follows an exponential distribution with a mean of one) will expect to live for 20.4 more years as given by equation (4.1).
According to Jensen’s Inequality, we can show that the average life expectancy is smaller than the life expectancy of an individual with an average vitality level:
[…]
The above result suggests that even if an individual can objectively determine their health level relative to the population, it does not guarantee an unbiased estimate of their survival probabilities. However, if the individual judges their vitality level based on quantiles rather than the average (e.g., the median of the population vitality level), they will be able to arrive at an unbiased estimate of their life expectancy.
@@ 4.4.1 Permanent disability model
Consider that there exists a threshold $\omega>0$ such that when the vitality falls below $\omega$, the person becomes disabled. Consider the specification with an exponentially distributed initial vitality (C1) and a Gompertz law specified depletion rate (C2). The resulting vitality model will form a permanent disability model, since $Y(t)$ is a strictly increasing function of time. The probability that a healthy person with initial vitality above $\omega$ will remain healthy for $T$ years is
$\Pr\left(V(0)-Y(T)>w\big{|}V(0)>w\right)$ $=\frac{\int_{0}^{\infty}e^{-v}\cdot\mathbb{1}_{v-Y(T)>\omega}\mathop{}\!\mathrm{d}v}{\int_{0}^{\infty}e^{-v}\cdot\mathbb{1}_{v>\omega}\mathop{}\!\mathrm{d}v}$
$=\frac{\int_{Y(T)+\omega}^{\infty}e^{-v}\mathop{}\!\mathrm{d}v}{\int_{w}^{\infty}e^{-v}\mathop{}\!\mathrm{d}v}=\exp(-Y(T)),$
which is equivalent to the survival probability, and the disability probability is simply $1-\exp(-Y(T))$. This result holds even if we assume that each individual has a different disability threshold, such that the threshold $\Omega$ has a density function $\pi(\omega)$. In this case, the probability of remaining healthy for $T$ years is
$\frac{\int_{0}^{\infty}e^{-v}\cdot\int_{0}^{\infty}\pi(\omega)\cdot\mathbb{1}_{v-\omega-Y(T)>0}\mathop{}\!\mathrm{d}\omega\mathop{}\!\mathrm{d}v}{\int_{0}^{\infty}e^{-v}\cdot\int_{0}^{\infty}\pi(\omega)\cdot\mathbb{1}_{v-\omega>0}\mathop{}\!\mathrm{d}\omega\mathop{}\!\mathrm{d}v}$ $=\frac{\int_{0}^{\infty}\pi(\omega)e^{-(\omega+Y(T))}\mathop{}\!\mathrm{d}\omega}{\int_{0}^{\infty}\pi(\omega)\cdot e^{-\omega}\mathop{}\!\mathrm{d}\omega}=\exp(-Y(T)).$
@@ 4.4.2 Disability model with recovery
For illustrative purposes, we consider the alternative Gompertz specification from Section 3.1.4 with a diffusion component (C3) modeled as a Brownian motion multiplied by a constant volatility $\sigma$, and set a constant threshold $\omega$ for disability. The $T$-year recovery probability for a disabled person aged $x$ is
$\frac{\int_{0}^{\omega}\Pr\left(v-Y(T)-W(T)>\omega,\text{ and }v-Y(t)-W(t)>0\text{ for any }t\in[0,T]\right)\mathop{}\!\mathrm{d}F_{0}(v)}{1-F_{0}(\omega)}.$
Since $Y(t)=\delta t$, the probability inside in the integral can be solved explicitly by setting $\widehat{W}(t)=W(t)+\frac{\delta}{\sigma}t$, which is a Brownian Motion with drift $\frac{\delta}{\sigma}$. Denote
$\widehat{M}(T)=\max_{0\leq t\leq T}\{W(t)+Y(t)\},$
the joint density of $(\widehat{M}(T),\widehat{W}(T))$ can be found as
$f(m,w)=\frac{2(2m-w)}{T\sqrt{2\pi T}}e^{\frac{\delta w}{\sigma}-\frac{\delta^{2}T}{2\sigma^{2}}-\frac{(2m-w)^{2}}{2T}},w\leq m,m\geq 0,$
and is zero otherwise (see Shreve et al., (2004) for a detailed derivation). Consequently, the probability inside the integral becomes
$\Pr\left(\widehat{M}(T)<\frac{v}{\sigma},\quad\widehat{W}(T)<\frac{v-\omega}{\sigma}\right)=\int_{0}^{v}\int_{0}^{v-\omega}f(m,w)\mathop{}\!\mathrm{d}m\mathop{}\!\mathrm{d}w.$
The recovery probability can then be determined by numerically solving the triple integral.

### v3

@@ 3.2.3 A vitality model with age-adjusted period-cohort effect
    The initial vitality $V_{x,y}(0)$ for an individual born in year $y$ and age $x$ at time $t=0$ follows
    $V_{x,y}(0)\sim\text{Exp}\left(\Gamma(y)^{\frac{x_{c}-x}{x_{c}}}\right),$
    where $x_{c}$ is a large and fixed value (e.g., $x_{c}=100$) and $\Gamma(y)$ is governed by a stochastic process (e.g., the geometric Brownian motion given by equation (10)).
@@
    The trend component $Y(t)$ is specified by $\mathop{}\!\mathrm{d}Y(t)=b(t)c(t)^{x+t}\mathop{}\!\mathrm{d}t$, $t\geq 0$, where the dynamics of $b(t)$ and $c(t)$ are the same as those presented in Sections 3.2.1 and 3.2.2.
The above specification provides flexibility in modeling the interactions between cohorts and ages. More specifically, $\Gamma(y)$ captures the general level of initial vitality at birth for cohort year $y$, with the term $\frac{x_{c}-x}{x_{c}}$ reflecting an age-adjusted cohort effect as the individual ages from 0 to $x_{c}$. When $x=0$, the initial vitality follows $\text{Exp}(\Gamma(y))$ and the specification coincides with the one presented in Section 3.2.2. When $x=x_{c}$, the initial vitality follows $\text{Exp}(1)$ and the specification coincides with the one presented in Section 3.2.1.
It is worth noting that, if $\Gamma(y)>1$, the cohort effect will diminish as the individual ages, which coincides with the cohort effect observed in the M8 model.
@@ 4 Model Applications
In this section, we provide a number of illustrative examples to showcase the application of a vitality mortality model in various research or practical problems. Specifically, we examine the following topics:
@@
We believe that the vitality mortality model differs from existing mortality models not only in terms of technical details, but also by offering new interpretations and insights that cannot be explained using existing mortality models.
@@ 4.1 Life insurance valuation
If the density function of $\tau(v)$ can be explicitly derived, then the life expectancy, annuity price, and insurance price can also be expressed explicitly in an integral form. This is the case for our proposed model if the diffusion component is not considered. When the density function of $\tau(v)$ cannot be explicitly written, an approximation method is then needed to provide accurate estimates for life expectancy and prices. For example, consider a vitality mortality model with initial vitality, trend, and diffusion (i.e., components C1-C3), then the problem becomes finding the first exit density of Brownian motion with a curved boundary. A number of studies, including Jennen and Lerche, (1981), Jennen, (1985), and Durbin and Williams, (1992), has provided accurate and easy-to-implement approximation methods to handle this problem. Skiadas and Skiadas, (2010) has specifically applied approximation methods to vitality modeling. For more details on approximating the density function of the death time when C4 is absent, interested readers are referred to Appendix E.
@@ 4.2 Optimal investment and consumption strategies
One promising application of vitality modeling is in lifetime planning, e.g., Merton’s portfolio problem. Individuals, when making decisions about their investment portfolio, consumption, insurance purchases, medical expenses, educational investment, etc., will take into account their health level at the time. The vitality mortality model provides a natural way to frame the optimization problem. There exist a few studies that utilize the vitality-based approach. For example, Chen et al., (2022) modeled vitality as a geometric Brownian motion to study the optimal insurance decision, and Cheng et al., (2023) modeled vitality using a jump process to study the optimal medical investment. In this subsection, we provide an illustration of Merton’s portfolio consumption problem under our vitality mortality model.
@@ 4.2.1 The market and vitality processes
Following the seminal work of Merton, (1969), we construct a continuous-time model for the financial market and apply the vitality model (7) without jumps. Denote $(\Omega,\mathbb{F},\mathbb{P})$ as a complete probability space with a right-continuous filtration generated by a two-dimensional standard Brownian Motion $(B_{S}(t),B_{V}(t))$. An individual allocates two representative assets, one risk-free $S_{0}(t)$ and one risky $S_{1}(t)$ with the following dynamics:
@@
$\mathop{}\!\mathrm{d}S_{1}(t)$ $=(r+\theta\sigma_{S})S_{1}(t)\mathop{}\!\mathrm{d}t+\sigma_{S}S_{1}(t)\mathop{}\!\mathrm{d}B_{S}(t),\quad S_{1}(0)>0,$
@@
Let $\pi(t)$ be the portfolio weight invested in the risky asset, and $\zeta(t)$ be the consumption amount. Then the asset $A(t)$ of an individual has the following dynamics
$\mathop{}\!\mathrm{d}A(t)=A(t)\left[r+\pi(t)\theta\sigma_{S}\right]\mathop{}\!\mathrm{d}t+\sigma_{S}\pi(t)A(t)\mathop{}\!\mathrm{d}B_{S}(t)-\zeta(t)\mathop{}\!\mathrm{d}t,\quad A(0)=A_{0}.$
We now adopt the vitality mortality model from Eq.(7) without jumps, which generalizes the Gompertz law if the initial vitality follows the Gompertz distribution; that is,
@@ 4.2.2 The optimal investment and consumption problem
where $\tau$ is the first passage time that vitality level crosses zero, $\beta>0$ is the time-preference rate, $u(\zeta)=\ln\zeta$ is the utility function (log-utility is selected for illustration), $\lambda\geq 0$ is the weight given to the utility of terminal asset (i.e., bequest), $\mathbb{E}_{t,a,v}=\mathbb{E}[\cdot|t,A(t)=a,V(t)=v]$, and $\Pi$ is the admissible set with definition given in Appendix C. Notice that the optimization is time-independent since $V(t)$ is stationary and thus the value function does not depend on the time $t$.
@@
As for comparison, we have Merton’s optimal portfolio and consumption strategies presented in Appendix C. The optimal investment strategy follows the same as Merton’s optimal portfolio in an infinite time horizon, and the consumption strategy also collapsed to Merton’s case (without mortality) if vitality is infinite.
@@ 4.2.3 Further remarks
It is not surprising to observe that the individual’s consumption depends on the vitality level, but what is interesting here is that they are not always negatively correlated, as one would expect. When the bequest motive is moderate or low (i.e., $\lambda<1/\beta$), a reasonable consumption profile is observed such that a higher vitality represents a longer time to live with, and thus a lower consumption must be made to make sure an individual will not run out of money before they are deceased.
However, when the bequest motive is significantly large (i.e., $\lambda>1/\beta$), decreasing the vitality actually increases the consumption, a counter-intuitive result at first glance. This is due to the fact that individuals are overly concerned about their bequest, and thus leads to an overly conservative consumption initially. Once they are aging and the time for death becomes more certain, they become more aware of their over-saving and may wish to increase their consumption a bit, as it would not jeopardize their bequest goals but will significantly enhance their lifetime utility. Figure 3 illustrates the sample paths for the optimal consumption rate (as a % of wealth) and the corresponding vitality process. The left panel represents a low bequest motive individual ($\lambda<1/\beta$), and the right panel represents a high bequest motive individual ($\lambda>1/\beta$). It is evident that a higher bequest motive results in a significantly lower consumption profile and a positive correlation between the consumption rate and vitality level. Conversely, when the bequest motive is low, the opposite pattern is observed.
Figure 3: Sample trajectories of vitality process (transparent lines) and optimal consumption rate (solid lines) for relatively low bequest motive (left) and high bequest motive (right).
@@ 4.3 Subjective beliefs
Another interesting observation that can be reflected by our vitality model is the distinction between subjective belief in health status and subjective belief in life expectancy. Specifically, if a person correctly self-assesses as having an average health level among the whole population, they may still over- or under-estimate their life expectancy if they falsely believe that their life expectancy will also be the average among the whole population. Using the specification from Section 3.1.1 with $b=0.0001744$ and $c=1.082$ as an illustrative example, the population average life expectancy at age 60 is 17 years. However, a 60-year-old individual who believes that they have an average initial vitality level (which is one since the initial vitality follows an exponential distribution with a mean of one) will expect to live for 20.4 more years as given by equation (4.1). This observation can be summarized in the following proposition.
Proposition 2.
Given that $V(0)\sim F_{0}$ where $F_{0}$ is a distribution function with positive support, the deterministic vitality depletion rate is strictly positive $\mu(t)>0$, and there are no diffusion and jump components, then
$\mathbb{E}[\tau]\leq\tau(\mathbb{E}[V(0)]),$
where $\tau(\cdot)$ is the conditional death time defined in Section 4. The average life expectancy is smaller than the life expectancy of an individual with an average vitality level.
The proof immediately follows from the fact that the inverse function of a strictly increasing function is strictly concave, followed by Jensen’s inequality. For example, for the vitality counterparts of traditional mortality law displayed in Section 3.1.1, $Y(t)$ is a strictly increasing function of time, then, the death time given the initial vitality value is a concave function. The proposition suggests that even if an individual can objectively determine their health level relative to the population, it does not guarantee an unbiased estimate of their survival probabilities. However, if the individual judges their vitality level based on quantiles rather than the average (e.g., the median of the population vitality level), they will be able to arrive at an unbiased estimate of their life expectancy.
@@ 4.4.1 Permanent disability model
Consider that there exists a threshold $\psi>0$ such that when the vitality falls below $\psi$, the person becomes disabled. Consider the specification with an exponentially distributed initial vitality (C1) and a Gompertz law specified depletion rate (C2). The resulting vitality model will form a permanent disability model, since $Y(t)$ is a strictly increasing function of time. The probability that a healthy person with initial vitality above $\psi$ will remain healthy for $T$ years is
$\Pr\left(V(0)-Y(T)>\psi\big{|}V(0)>\psi\right)$ $=\frac{\int_{0}^{\infty}e^{-v}\cdot\mathbb{1}_{v-Y(T)>\psi}\mathop{}\!\mathrm{d}v}{\int_{0}^{\infty}e^{-v}\cdot\mathbb{1}_{v>\psi}\mathop{}\!\mathrm{d}v}$
$=\frac{\int_{Y(T)+\psi}^{\infty}e^{-v}\mathop{}\!\mathrm{d}v}{\int_{\psi}^{\infty}e^{-v}\mathop{}\!\mathrm{d}v}=\exp(-Y(T)),$
which is equivalent to the survival probability, and the disability probability is simply $1-\exp(-Y(T))$. This result holds even if we assume that each individual has a different disability threshold, such that the threshold $\Psi$ has a density function $\pi(\psi)$. In this case, the probability of remaining healthy for $T$ years is
$\frac{\int_{0}^{\infty}e^{-v}\cdot\int_{0}^{\infty}\pi(\psi)\cdot\mathbb{1}_{v-\psi-Y(T)>0}\mathop{}\!\mathrm{d}\psi\mathop{}\!\mathrm{d}v}{\int_{0}^{\infty}e^{-v}\cdot\int_{0}^{\infty}\pi(\psi)\cdot\mathbb{1}_{v-\psi>0}\mathop{}\!\mathrm{d}\psi\mathop{}\!\mathrm{d}v}$ $=\frac{\int_{0}^{\infty}\pi(\psi)e^{-(\psi+Y(T))}\mathop{}\!\mathrm{d}\psi}{\int_{0}^{\infty}\pi(\psi)\cdot e^{-\psi}\mathop{}\!\mathrm{d}\psi}=\exp(-Y(T)).$
@@ 4.4.2 Disability model with recovery
For illustrative purposes, here we consider the simplest model that is able to capture the recovery probability, specifically, we consider the alternative Gompertz specification with a constant depletion rate and a diffusion component (C3) modeled as a Brownian motion multiplied by a constant volatility, along with a constant disability threshold. The result is summarized in the following proposition.
Proposition 3.
Consider the vitality model with an initial vitality level $V(0)\sim F_{0}$, a constant vitality depletion rate $\mu(t)=\delta$, and a diffusion component as a Brownian motion with volatility $\sigma$. Consider a constant disability threshold $\psi$, then the $T$-year disabled probability for a healthy person aged $x$ is
$\int_{\psi}^{\infty}\frac{1}{1-F_{0}(\psi)}\cdot\left(\int_{0}^{\frac{v}{\sigma}}\int_{\frac{v-\psi}{\sigma}}^{\frac{v}{\sigma}}f(m,w;T)\mathop{}\!\mathrm{d}w\mathop{}\!\mathrm{d}m\right)\mathop{}\!\mathrm{d}F_{0}(v),$
and the recovery probability for a disabled person aged $x$ is
$\int_{0}^{\psi}\frac{1}{F_{0}(\psi)}\cdot\left(\int_{0}^{\frac{v}{\sigma}}\int_{0}^{\frac{v-\psi}{\sigma}}f(m,w;T)\mathop{}\!\mathrm{d}w\mathop{}\!\mathrm{d}m\right)\mathop{}\!\mathrm{d}F_{0}(v),$
where
$f(m,w;T)=\begin{cases}\frac{2(2m-w)}{T\sqrt{2\pi T}}e^{\frac{\delta w}{\sigma}-\frac{\delta^{2}T}{2\sigma^{2}}-\frac{(2m-w)^{2}}{2T}},&w\leq m,m\geq 0\\ \\ 0,&\text{elsewhere}.\end{cases}$
The proof is given in Appendix D. The probabilities can be evaluated by numerically solving the triple-integral. Notice that the choice of the disability threshold here does not impact the death probability, suggesting an approach to seamlessly integrate disability modeling and mortality modeling, which contrasts with traditional multi-state modeling for disability. Of course, the model above might be overly simplified; further exploration in model extension and empirical examination with disability data would be necessary.